in such a way that its height, h meters after t seconds is given by the equation h.-49%28 7t +42 How Atoy rocket i launched froma 42 high plat ong will it take for the rocket to hit the ground? The toy rocket will hit the ground after approximately □ seconds Type an integer or a decimal Round to the nearest hundredth as needed)
Answer #1

Height is given as:

h(t) = -4.9t^2+28.7t +4.2

So when rocket will hnit the ground it's height is equal to zero.

So we can write

-4.9t^2+28.7t - 4.2 = 0

0.7t^2-4.1t - 0.6 = 0

0.7t^2-4.2t + 0.1t - 0.6 = 0

0.7t(t-6)+ 0.1(t - 6) = 0

(t-6)( 0.7t +0.1)= 0

So t - 6 = 0 or 0.7t +0.1 = 0

So t = 6 or t = -\frac{1}{7}

But t can't be negative.

So t = 6

So we can say 6 sec rocket will hit the ground.